# Heat current and fluctuations between a dissipative qubit and a monitor under continuous measurement and feedback

### v2

@@ 3.3 Zero heat current
Here, $\Delta N_{t}$ is the Poisson increment and satisfies $\Delta N_{t}\Delta N_{t}=\Delta N_{t}$. It takes $\Delta N_{t}=1$ when the monitor detects that the qubit is in |$\ket{m}$ at time $t$ and $\Delta N_{t}=0$ otherwise. The probability of $\Delta N_{t}=1$ is given by $\mathbb{E}[\Delta N_{t}]=\gamma\bra{m}\rho_{\rm c}(t)\ket{m}\Delta t$, where $\mathbb{E}[\cdots]$ denotes the ensemble average over the measurement outcomes. The unraveled equation (3.3) reproduces the Lindblad equation (4) after the ensemble average, $\rho(t)=\mathbb{E}[\rho_{\rm c}(t)]$. For $\Delta N_{t}=1$, the conditional reduced density matrix jumps to the feedback state, $\rho_{\rm c}(t+\Delta t)=\ket{n}\bra{n}$ |, called a quantum jump, and then the ensemble-averaged qubit energy at $t+\Delta t$ is $(\gamma\Delta/2)\rho_{mm}{\rm tr}_{0}[H_{0}P_{n}]\Delta t$. Since this energy change by the quantum jump is due to the quantum measurement and feedback, the heat current induced by the monitor is $J^{(2)}=\gamma\rho_{mm}{\rm tr}_{0}[H_{0}P_{n}]-\gamma\rho_{mm}{\rm tr}_{0}[H_{0}\rho]$. For $\Delta N_{t}=0$, since the reduced density matrix evolves as $\dot{\rho}=-i[H_{0},\rho]+\mathcal{D}_{\rm B}[\rho]+\mathcal{D}_{\rm M}^{(1)}[\rho]$, the heat current by the monitor is $J^{(1)}={\rm tr}_{0}[H_{0}\mathcal{D}_{\rm M}^{(1)}[\rho]]=\gamma\rho_{mm}{\rm tr}_{0}[H_{0}\rho]-(\gamma/2){\rm tr}_{0}[H_{0}\{P_{m},\rho\}]$ [46]. Note that $J=J^{(1)}+J^{(2)}$. Therefore, the first term of the heat current (6) represents the backaction of the detection ($\Delta N=1$) and the second term the backaction of no detection ($\Delta N=0$). The heat current does not flow between the qubit and the monitor when they are balanced, i.e., $J^{(1)}+J^{(2)}=0$, corresponding to $\rho_{mm}(\cos\theta_{m}-\cos\theta_{n})=\sin\theta_{m}{\rm Re}[\rho_{m\bar{m}}]$. One can confirm that it reproduces $\theta_{m}=0$ and $\pi$ for the measurement-only case ($\theta_{m}=\theta_{n}$).
@@ 3.4 Quantum measurement cooling
Figure 3: (a) Heat current flowing out of the cold heat bath to the qubit $J_{c}$ as a function of $\theta_{m}$ and $\theta_{n}$, obtained by solving the quantum master equation (4), for $T_{h}/\Delta=1.5$, $T_{c}/\Delta=1$, $\alpha_{h}=\alpha_{c}=0.01$, and $\gamma/\Delta=0.01$. The quantum measurement cooling occurs in the red region ($J_{c}>0$). (b) The COP as a function of $J_{c}$ ($>0$) at symmetric point $\theta_{m}+\theta_{n}=\pi$ for different temperature biases.
The continuous measurement and feedback can induce the qubit cooling ($J<0$), as discussed in Sec. 3.1. This indicates the possibility of the quantum measurement cooling [22], which is the energy extraction from the colder heat bath by quantum measurement. Now, we consider the two heat baths ($r=h,c$) with different temperatures ($T_{h}<T_{c}$). The sign of the heat current flowing out of the cold heat bath determines if the quantum measurement cooling occurs. The heat current flowing out of the cold and hot heat baths to the qubit are defined as
@@ 4 Fluctuation
So far, we considered the steady-state heat current after taking the ensemble average over the measurement outcomes. However, since the measurements are a stochastic process, the heat current fluctuates around the averaged heat current even at the steady-state limit. To characterize the fluctuation of the heat current, we introduce the power spectrum as
@@
where $C_{J}(t,t^{\prime};t_{0})=\mathbb{E}[\delta J_{\rm c}(t+t_{0})\delta J_{\rm c}(t^{\prime}+t_{0})]$ is the correlation function of the heat current fluctuation around the averaged heat current, $\delta J_{\rm c}(t)\equiv J_{\rm c}(t)-J(t)$. The conditional heat current $J_{\rm c}(t)$ is decomposed into the contributions from the backaction of no detection ($\Delta N_{t}=0$) and the quantum jump ($\Delta N_{t}=1$),
@@
respectively. Here, $t_{0}$ is large enough that the qubit reaches the steady state and the time $\mathcal{T}$ is sufficiently long for the current correlation to vanish and will eventually be taken to infinity. Below, we address the correlation function of the heat change during the infinitesimal time $\Delta t$,
@@
where the conditional heat change is $Q_{\rm c}(t)=Q_{\rm c}^{(1)}(t)+Q_{\rm c}^{(2)}(t)=J_{\rm c}^{(1)}(t)\Delta t+J_{\rm c}^{(2)}(t)\Delta t$.
@@ 4.1 Poisson noise
The frequency-independent term (20a) of the power spectrum comes from the local correlation in time, $C(0)=\mathbb{E}\left[Q_{\rm c}^{2}\right]-\mathbb{E}[Q_{\rm c}]^{2}$. In the following, we drop the time index for the steady state. At $\Delta t\to 0$, the correlation of the heat changes by the quantum jump process, $\mathbb{E}[(Q_{\rm c}^{(2)})^{2}]$ is predominant. Therefore, the frequency-independent power spectrum is obtained as
@@
where $Q_{\rm jump}={\rm tr}_{0}[H_{0}(P_{n}-\rho)]$ and $J^{(2)}=\mathbb{E}[Q_{\rm c}^{(2)}]/\Delta t=\gamma\rho_{mm}{\rm tr}_{0}[H_{0}(P_{n}-\rho)]$ are the heat change and the steady-state heat current induced by the quantum jump process. This expression corresponds to the Poisson noise for non-interacting electrons in a one-dimensional nanowire with weak transmission; $S_{\rm Poisson}(t)\approx 2eI(t)$, where $I(t)$ represents the mean electronic current [31, 32]. In this work, we call the frequency-independent contribution of the power spectrum $S_{0}$ as the “Poisson” noise. In the weak coupling regime, the Poisson noise is approximated as
@@
Figure 4 (a) shows the Poisson noise for $\Gamma_{+}/\Delta=0.3$, $\Gamma_{-}/\Delta=0.15$, and $\gamma/\Delta=0.01$. The Poisson noise varies, depending on the measurement and feedback states, and vanishes, as shown in the solid line, when the qubit energy of the steady state matches that of the feedback state, ${\rm tr}_{0}[H_{0}\rho]={\rm tr}_{0}[H_{0}P_{n}]$. The line of $S_{0}=0$ is almost independent of the measurement state. The heat change due to the quantum jump is expressed as $Q_{\rm jump}=-(\Delta/2)(\cos\theta_{n}+\braket{\sigma_{z}})$, and the $\gamma$ dependence of $\braket{\sigma_{z}}$ is small for $\gamma\ll\gamma_{\pm}$, i.e., it is primarily determined by the dissipation due to the heat baths. Thus, $Q_{\rm jump}$ is almost independent of the measurement state, while it strongly depends on the feedback state. Since $\braket{\sigma_{z}}<0$ and $\braket{\sigma_{z}}$ is almost independent of $\gamma$, equivalently $\theta_{m}$ and $\theta_{n}$, $|Q_{\rm jump}|$ takes maximum at $\theta_{n}=\pi$ and its value is positive, $Q_{\rm jump}>0$. Concerning another contribution of the Poisson noise, $\rho_{mm}$, because $\rho_{mm}\approx(1-\braket{\sigma_{z}}\cos\theta_{m})/2$ for $\gamma\ll\gamma_{\pm}$, it is almost independent of the feedback state and reaches maximum at $\theta_{m}=0$. Therefore, the Poisson noise has a maximum value at $\theta_{m}=0$ and $\theta_{n}=\pi$, as shown in Fig. 4. There, more heat is transferred between the qubit and the monitor more frequently.
We also show the $\theta_{m}$ dependence of the Poisson noise for the measurement-only protocol, |$\ket{m}=\ket{n}$, in Fig. 4 (b). The expression for the Poisson noise (22), describes the numerical simulation well. The Poisson noise is not symmetric with respect to $\theta_{m}=\pi/2$, unlike the steady-state heat current. This is reflected in the characteristic statistics induced by the continuous measurement and feedback, which differs from the case of the electronic current. The resulting fluctuations provide direct access to the quantum jump, which is indistinguishable from the backaction of no detection at the level of the heat current. The characteristic statistics arise from the fact the probability of a measurement occurring depends on the qubit state, and the heat current is also affected by the measurement backaction, discussed in Sec. 4.2.
@@ 4.2 Backaction noise
After the quantum jump, heat exchanges between the qubit and the monitor to reach the steady state as a measurement backaction. The heat current fluctuation induced by the measurement backaction is characterized by the non-local correlation in time, $C_{1}(t)$, providing the frequency-dependent contribution of the power spectrum $S_{1}(\omega)$, called the backaction noise. At the dc limit ($\omega=0$), the power spectrum is written as
@@
We plot the backaction noise in the plane of ($\theta_{m},\theta_{n}$) in Fig. 5 (a) for the same parameters of the Poisson noise in Fig. 4. The magnitude of the backaction noise is smaller than the Poisson noise by one order, as the excess heat is on the order of $\Delta\gamma/\gamma_{+}$ and $Q_{\rm jump}$ is on the order of $\Delta$. The backaction noise is negative in a large range of ($\theta_{m}$, $\theta_{n}$). The negative backaction noise indicates that, from Eq. (27), the heat change by the quantum jump and the excess heat have opposite signs. This is the measurement backaction; when the qubit energy is stored from the monitor by the detection ($Q_{\rm jump}>0$), less heat flows out of the monitor to the qubit in the dynamics after the detection compared with the steady-state heat current ($Q_{\rm ex}<0$), and vice versa.
@@
The backaction noise has a minimum value around $\theta_{m}=0$ and $\theta_{n}=\pi$. The excess heat is the difference between the total energy transferred from the monitor to the qubit until the post-selected state reaches the steady state and the steady-state heat current during that time. Roughly speaking, $|Q_{\rm ex}|$ becomes larger as $J(t=0;0)=-(\gamma\Delta/4)\sin(\theta_{m}-\theta_{n})\sin\theta_{n}$ and $J$ are more different because, for $\gamma\ll\gamma_{\pm}$, the relaxation rate to the steady state heat current is dominantly determined by the heat baths, and the measurement effect to the rate is small. The oscillation of the transient heat current due to the quantum coherence affects the excess heat, but it reduces $|Q_{\rm ex}|$ because the integral of the product of a trigonometric function and an exponential decay function becomes smaller than that without oscillation when the oscillation frequency, $\sim\Delta$ is larger than the relaxation rate, $\sim\gamma_{+}$. At $\theta_{m}=0$ and $\theta_{n}=\pi$, $J$ ($>0$) takes maximum as discussed in Sec. 3 and the heat current relaxes to the steady-state one without the oscillation while $J(t=0;0)=0$. Then, $|Q_{\rm ex}|$ is maximum and $Q_{\rm ex}$ is negative around $\theta_{m}=0$ and $\theta_{n}=\pi$. Hence, from the fact that $\rho_{mm}Q_{\rm jump}$ is positive and its absolute value has a maximum value at $\theta_{m}=0$ and $\theta_{n}=\pi$, as discussed in Sec. 4.1, the backaction noise is negative and its absolute value reaches maximum around $\theta_{m}=0$ and $\theta_{n}=\pi$. In contrast, when $\theta_{m}=\theta_{n}=0$ or $\pi$, the excess heat vanishes, resulting in $S_{1}=0$, because the measurement commutes with the qubit Hamiltonian. In addition, the backaction noise takes 0 around $\theta_{n}/\pi\approx 0.4$, corresponding to $Q_{\rm ex}=0$ and $Q_{\rm jump}+Q_{\rm ex}/2=0$, which will be discussed below.
@@
Figure 6: Heat change due to the quantum jump, $Q_{\rm jump}$ (solid line), and excess heat, $Q_{\rm ex}$ (dashed line), as a function of $\theta_{m}$ for the measurement-only case ($\theta_{m}=\theta_{n}$). For a large range of $\theta_{m}$, $Q_{\rm jump}$ and $Q_{\rm ex}$ have opposite signs, but they are both negative when they are crossing 0. The inset shows an enlarged view where both $Q_{\rm jump}$ and $Q_{\rm ex}$ are negative near $\theta_{m}/\pi\approx 0.39$. The difference between $Q_{\rm jump}$ and $Q_{\rm jump}+Q_{\rm ex}/2$ is not visible in this plot.
@@
As mentioned above, the backaction noise has negative values for a large range of ($\theta_{m},\theta_{n}$), but it shows the small positive value, the order of $10^{-6}$, at $\theta_{m}/\pi\approx 0.3-0.5$ and $\theta_{n}/\pi\approx 0.4$ near $S_{0}=0$. It, here, violates $Q_{\rm jump}Q_{\rm ex}<0$. This is attributed that the energy loss or gain of the qubit due to quantum measurement can also be compensated by the heat baths during the transient dynamics. Therefore, when focusing on the heat exchange between the monitor and the qubit, it is possible to slightly violate $Q_{\rm jump}Q_{\rm ex}<0$ around $Q_{\rm jump},Q_{\rm ex}\approx 0$. We plot $Q_{\rm jump}$ and $Q_{\rm ex}$ for the measurement-only case in Fig. 6, and it is indeed observed that both values are negative within the small range of $\theta_{m}/\pi\approx 0.381-0.393$.

### v3

@@ 3.3 Zero energy exchange
Here, $\Delta N_{t}$ is the Poisson increment and satisfies $\Delta N_{t}\Delta N_{t}=\Delta N_{t}$. It takes $\Delta N_{t}=1$ when the monitor detects that the qubit is in |$\ket{m}$ at time $t$ and $\Delta N_{t}=0$ otherwise. The probability of $\Delta N_{t}=1$ is given by $\mathbb{E}[\Delta N_{t}]=\gamma\bra{m}\rho_{\rm c}(t)\ket{m}\Delta t$, where $\mathbb{E}[\cdots]$ denotes the ensemble average over the measurement outcomes. The unraveled equation (3.3) reproduces the Lindblad equation (4) after the ensemble average, $\rho(t)=\mathbb{E}[\rho_{\rm c}(t)]$. For $\Delta N_{t}=1$, the conditional reduced density matrix jumps to the feedback state, $\rho_{\rm c}(t+\Delta t)=\ket{n}\bra{n}$ |, called a quantum jump, and then the ensemble-averaged qubit energy at $t+\Delta t$ is $(\gamma\Delta/2)\rho_{mm}{\rm tr}_{0}[H_{0}P_{n}]\Delta t$. Since this energy change by the quantum jump is due to the quantum measurement and feedback, the energy flow induced by the monitor is $J^{(2)}=\gamma\rho_{mm}{\rm tr}_{0}[H_{0}P_{n}]-\gamma\rho_{mm}{\rm tr}_{0}[H_{0}\rho]$. For $\Delta N_{t}=0$, since the reduced density matrix evolves as $\dot{\rho}=-i[H_{0},\rho]+\mathcal{D}_{\rm B}[\rho]+\mathcal{D}_{\rm M}^{(1)}[\rho]$, the energy flow by the monitor is $J^{(1)}={\rm tr}_{0}[H_{0}\mathcal{D}_{\rm M}^{(1)}[\rho]]=\gamma\rho_{mm}{\rm tr}_{0}[H_{0}\rho]-(\gamma/2){\rm tr}_{0}[H_{0}\{P_{m},\rho\}]$ [46]. Note that $J=J^{(1)}+J^{(2)}$. Therefore, the first term of the energy flow (6) represents the backaction of the detection ($\Delta N=1$) and the second term the backaction of no detection ($\Delta N=0$). The energy exchange does not occur between the qubit and the monitor when they are balanced, i.e., $J^{(1)}+J^{(2)}=0$, corresponding to $\rho_{mm}(\cos\theta_{m}-\cos\theta_{n})=\sin\theta_{m}{\rm Re}[\rho_{m\bar{m}}]$. One can confirm that it reproduces $\theta_{m}=0$ and $\pi$ for the measurement-only case ($\theta_{m}=\theta_{n}$).
@@ 3.4 Quantum measurement cooling
Figure 3: (a) Heat current from the cold heat bath to the qubit $J_{c}$ as a function of $\theta_{m}$ and $\theta_{n}$, obtained by solving the quantum master equation (4), for $T_{h}/\Delta=1.5$, $T_{c}/\Delta=1$, $\alpha_{h}=\alpha_{c}=0.01$, and $\gamma/\Delta=0.01$. The quantum measurement cooling occurs in the red region ($J_{c}>0$). (b) The COP as a function of $J_{c}$ ($>0$) at symmetric point $\theta_{m}+\theta_{n}=\pi$ for different temperature biases.
The continuous measurement and feedback can induce the qubit cooling ($J<0$), as discussed in Sec. 3.1. This indicates the possibility of the quantum measurement cooling [22], which is the energy extraction from the colder heat bath by quantum measurement. Now, we consider the two heat baths ($r=h,c$) with different temperatures ($T_{h}<T_{c}$). The sign of the heat current from the cold heat bath determines if the quantum measurement cooling occurs. The heat currents from the cold and hot heat baths to the qubit are defined as
@@ 4 Fluctuation
So far, we considered the steady-state energy flow after taking the ensemble average over the measurement outcomes. However, since the measurements are a stochastic process, the energy flow fluctuates around the averaged energy flow even at the steady-state limit. To characterize the fluctuation of the energy flow, we introduce the power spectrum as
@@
where $C_{J}(t,t^{\prime};t_{0})=\mathbb{E}[\delta J_{\rm c}(t+t_{0})\delta J_{\rm c}(t^{\prime}+t_{0})]$ is the correlation function of the energy flow fluctuation around the averaged energy flow, $\delta J_{\rm c}(t)\equiv J_{\rm c}(t)-J(t)$. The conditional energy flow $J_{\rm c}(t)$ is decomposed into the contributions from the backaction of no detection ($\Delta N_{t}=0$) and the quantum jump ($\Delta N_{t}=1$),
@@
respectively. Here, $t_{0}$ is large enough that the qubit reaches the steady state and the time $\mathcal{T}$ is sufficiently long for the current correlation to vanish and will eventually be taken to infinity. Below, we address the correlation function of the energy change during the infinitesimal time $\Delta t$,
@@
where the conditional energy change is $Q_{\rm c}(t)=Q_{\rm c}^{(1)}(t)+Q_{\rm c}^{(2)}(t)=J_{\rm c}^{(1)}(t)\Delta t+J_{\rm c}^{(2)}(t)\Delta t$.
@@ 4.1 Poisson noise
The frequency-independent term (20a) of the power spectrum comes from the local correlation in time, $C(0)=\mathbb{E}\left[Q_{\rm c}^{2}\right]-\mathbb{E}[Q_{\rm c}]^{2}$. In the following, we drop the time index for the steady state. At $\Delta t\to 0$, the correlation of the energy changes by the quantum jump process, $\mathbb{E}[(Q_{\rm c}^{(2)})^{2}]$ is predominant. Therefore, the frequency-independent power spectrum is obtained as
@@
where $Q_{\rm jump}={\rm tr}_{0}[H_{0}(P_{n}-\rho)]$ and $J^{(2)}=\mathbb{E}[Q_{\rm c}^{(2)}]/\Delta t=\gamma\rho_{mm}{\rm tr}_{0}[H_{0}(P_{n}-\rho)]$ are the energy change and the steady-state energy flow induced by the quantum jump process. This expression corresponds to the Poisson noise for non-interacting electrons in a one-dimensional nanowire with weak transmission; $S_{\rm Poisson}(t)\approx 2eI(t)$, where $I(t)$ represents the mean electronic current [31, 32]. In this work, we call the frequency-independent contribution of the power spectrum $S_{0}$ as the “Poisson” noise. In the weak coupling regime, the Poisson noise is approximated as
@@
Figure 4 (a) shows the Poisson noise for $\Gamma_{+}/\Delta=0.3$, $\Gamma_{-}/\Delta=0.15$, and $\gamma/\Delta=0.01$. The Poisson noise varies, depending on the measurement and feedback states, and vanishes, as shown in the solid line, when the qubit energy of the steady state matches that of the feedback state, ${\rm tr}_{0}[H_{0}\rho]={\rm tr}_{0}[H_{0}P_{n}]$. The line of $S_{0}=0$ is almost independent of the measurement state. The energy change due to the quantum jump is expressed as $Q_{\rm jump}=-(\Delta/2)(\cos\theta_{n}+\braket{\sigma_{z}})$, and the $\gamma$ dependence of $\braket{\sigma_{z}}$ is small for $\gamma\ll\gamma_{\pm}$, i.e., it is primarily determined by the dissipation due to the heat baths. Thus, $Q_{\rm jump}$ is almost independent of the measurement state, while it strongly depends on the feedback state. Since $\braket{\sigma_{z}}<0$ and $\braket{\sigma_{z}}$ is almost independent of $\gamma$, equivalently $\theta_{m}$ and $\theta_{n}$, $|Q_{\rm jump}|$ takes maximum at $\theta_{n}=\pi$ and its value is positive, $Q_{\rm jump}>0$. Concerning another contribution of the Poisson noise, $\rho_{mm}$, because $\rho_{mm}\approx(1-\braket{\sigma_{z}}\cos\theta_{m})/2$ for $\gamma\ll\gamma_{\pm}$, it is almost independent of the feedback state and reaches maximum at $\theta_{m}=0$. Therefore, the Poisson noise has a maximum value at $\theta_{m}=0$ and $\theta_{n}=\pi$, as shown in Fig. 4. There, more energy is transferred between the qubit and the monitor more frequently.
We also show the $\theta_{m}$ dependence of the Poisson noise for the measurement-only protocol, |$\ket{m}=\ket{n}$, in Fig. 4 (b). The expression for the Poisson noise (22), describes the numerical simulation well. The Poisson noise is not symmetric with respect to $\theta_{m}=\pi/2$, unlike the steady-state energy flow. This is reflected in the characteristic statistics induced by the continuous measurement and feedback, which differs from the case of the electronic current. The resulting fluctuations provide direct access to the quantum jump, which is indistinguishable from the backaction of no detection at the level of the energy exchange. The characteristic statistics arise from the fact the probability of a measurement occurring depends on the qubit state, and the energy exchange is also affected by the measurement backaction, discussed in Sec. 4.2.
@@ 4.2 Backaction noise
After the quantum jump, energy exchanges between the qubit and the monitor to reach the steady state as a measurement backaction. The energy flow fluctuation induced by the measurement backaction is characterized by the non-local correlation in time, $C_{1}(t)$, providing the frequency-dependent contribution of the power spectrum $S_{1}(\omega)$, called the backaction noise. At the dc limit ($\omega=0$), the power spectrum is written as
@@
We plot the backaction noise in the plane of ($\theta_{m},\theta_{n}$) in Fig. 5 (a) for the same parameters of the Poisson noise in Fig. 4. The magnitude of the backaction noise is smaller than the Poisson noise by one order, as the excess energy is on the order of $\Delta\gamma/\gamma_{+}$ and $Q_{\rm jump}$ is on the order of $\Delta$. The backaction noise is negative in a large range of ($\theta_{m}$, $\theta_{n}$). The negative backaction noise indicates that, from Eq. (27), the energy change by the quantum jump and the excess energy have opposite signs. This is the measurement backaction; when the qubit energy is stored from the monitor by the detection ($Q_{\rm jump}>0$), less energy flows out of the monitor to the qubit in the dynamics after the detection compared with the steady-state energy flow ($Q_{\rm ex}<0$), and vice versa.
@@
The backaction noise has a minimum value around $\theta_{m}=0$ and $\theta_{n}=\pi$. The excess energy is the difference between the total energy transferred from the monitor to the qubit until the post-selected state reaches the steady state and the steady-state energy flow during that time. Roughly speaking, $|Q_{\rm ex}|$ becomes larger as $J(t=0;0)=-(\gamma\Delta/4)\sin(\theta_{m}-\theta_{n})\sin\theta_{n}$ and $J$ are more different because, for $\gamma\ll\gamma_{\pm}$, the relaxation rate to the steady-state energy flow is dominantly determined by the heat baths, and the measurement effect to the rate is small. The oscillation of the transient energy flow due to the quantum coherence affects the excess energy, but it reduces $|Q_{\rm ex}|$ because the integral of the product of a trigonometric function and an exponential decay function becomes smaller than that without oscillation when the oscillation frequency, $\sim\Delta$ is larger than the relaxation rate, $\sim\gamma_{+}$. At $\theta_{m}=0$ and $\theta_{n}=\pi$, $J$ ($>0$) takes maximum as discussed in Sec. 3 and the energy flow relaxes to the steady-state one without the oscillation while $J(t=0;0)=0$. Then, $|Q_{\rm ex}|$ is maximum and $Q_{\rm ex}$ is negative around $\theta_{m}=0$ and $\theta_{n}=\pi$. Hence, from the fact that $\rho_{mm}Q_{\rm jump}$ is positive and its absolute value has a maximum value at $\theta_{m}=0$ and $\theta_{n}=\pi$, as discussed in Sec. 4.1, the backaction noise is negative and its absolute value reaches maximum around $\theta_{m}=0$ and $\theta_{n}=\pi$. In contrast, when $\theta_{m}=\theta_{n}=0$ or $\pi$, the excess energy vanishes, resulting in $S_{1}=0$, because the measurement commutes with the qubit Hamiltonian. In addition, the backaction noise takes 0 around $\theta_{n}/\pi\approx 0.4$, corresponding to $Q_{\rm ex}=0$ and $Q_{\rm jump}+Q_{\rm ex}/2=0$, which will be discussed below.
@@
Figure 6: Energy change due to the quantum jump, $Q_{\rm jump}$ (solid line), and excess energy, $Q_{\rm ex}$ (dashed line), as a function of $\theta_{m}$ for the measurement-only case ($\theta_{m}=\theta_{n}$). The parameters are the same as Fig. 4. For a large range of $\theta_{m}$, $Q_{\rm jump}$ and $Q_{\rm ex}$ have opposite signs, but they are both negative when they are crossing 0. The inset shows an enlarged view where both $Q_{\rm jump}$ and $Q_{\rm ex}$ are negative near $\theta_{m}/\pi\approx 0.39$. The difference between $Q_{\rm jump}$ and $Q_{\rm jump}+Q_{\rm ex}/2$ is not visible in this plot.
@@
As mentioned above, the backaction noise has negative values for a large range of ($\theta_{m},\theta_{n}$), but it shows the small positive value, the order of $10^{-6}$, at $\theta_{m}/\pi\approx 0.3-0.5$ and $\theta_{n}/\pi\approx 0.4$ near $S_{0}=0$. It, here, violates $Q_{\rm jump}Q_{\rm ex}<0$. This is attributed that the energy loss or gain of the qubit due to quantum measurement can also be compensated by the heat baths during the transient dynamics. Therefore, when focusing on the energy exchange between the monitor and the qubit, it is possible to slightly violate $Q_{\rm jump}Q_{\rm ex}<0$ around $Q_{\rm jump},Q_{\rm ex}\approx 0$. We plot $Q_{\rm jump}$ and $Q_{\rm ex}$ for the measurement-only case in Fig. 6, and it is indeed observed that both values are negative within the small range of $\theta_{m}/\pi\approx 0.381-0.393$.